### v1

# hep-ph/***FTUAM-21-$x$ IFT-UAM/CSIC-21-96

Abstract
The Kerr-Schild gauge is generalized to the case that the vector generating the deformation is not null. Contrary to naive expectations, this vector generates a finite expansion for the curvature tensor. We prove a theorem on the conditions for the deformed metric being Ricci flat, namely that the deformation vector must be geodesic of the background spacetime.
1 Introduction
One of the main problems in General Relativity is the non-linearity of the field equations, which means that the linear combination of solutions is not necessarily a solution, even at the free level. Given two Ricci flat metrics
$R_{\alpha\beta}[g^{(1)}_{\mu\nu}]=R_{\alpha\beta}[g^{(2)}_{\mu\nu}]=0$ (1)
the sum is not Ricci flat anymore
$R_{\alpha\beta}[g^{(1)}_{\mu\nu}+g^{(2)}_{\mu\nu}]\neq 0$ (2)
This non-linearity can be traced back the nonlinearity of the connection
$\Gamma^{\lambda}_{\alpha\beta}\left[g^{(1)}_{\mu\nu}+g^{(2)}_{\mu\nu}\right]\neq\Gamma^{\lambda}_{\alpha\beta}\left[g^{(1)}_{\mu\nu}\right]+\Gamma^{\lambda}_{\alpha\beta}\left[g^{(2)}_{\mu\nu}\right$] (3)
It is true that even if the connection were additive, the Ricci tensor would not share this property; but it seems a neccessary step in the good direction. The main technical problem here is the not additivity of the inverse of a sum of matrices
$(g^{(1)}_{\mu\nu}+g^{(2)}_{\mu\nu})^{-1}\neq(g^{(1)}_{\mu\nu})^{-1}+(g^{(2)}_{\mu\nu})^{-1}$ (4)
Let us examine this fact in more detail trying to discover exceptions to it; that is, situations where the expansion of the inverse matrix terminates. To be specific, when doing perturbation theory around a background gravitational field $\bar{g}_{\mu\nu}$
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\kappa h_{\mu\nu}$ (5)
the first step in the computation is to calculate the inverse metric, $g^{\mu\nu}$. This is only possible as an infinite power series, namely
$g^{\mu\nu}=\bar{g}^{\mu\nu}-\kappa h^{\mu\nu}+\kappa^{2}h_{\penalty 10000\ \tau}^{\mu}h^{\tau\nu}+\mathcal{O}(h^{2})$ (6)
let us note that indices are raised with the $\bar{g}_{\mu\nu}$ metric, and there is no closed formula in the general case. The same thing happens with the determinant of the metric.
A clever way out is the Kerr-Schild gauge [1, 2, 3, 4]
$g^{\mu\nu}=\bar{g}^{\mu\nu}-\kappa h^{\mu\nu}$ (7)
where the deformation obeys
$h_{\mu\lambda}h^{\lambda\nu}=0$
$h^{\lambda}_{\lambda}=0$ (8)
These equations imply that
$h_{\alpha\beta}=l_{\alpha}l_{\beta}$ (9)
with $l$ a null vector, $l^{2}=0$. In this case the Ricci tensor can be written as
$R_{\mu\nu}=\bar{R}_{\mu\nu}+R_{\mu\nu}^{\text{\tiny{(1)}}}+R_{\mu\nu}^{\text{\tiny{(2)}}}+R_{\mu\nu}^{\text{\tiny{(3)}}}$ (10)
where
$R_{\mu\nu}^{\text{\tiny{(1)}}}=\frac{\kappa}{2}\left[\bar{\nabla}_{\lambda}\bar{\nabla}_{\mu}h_{\nu}^{\penalty 10000\ \lambda}+\bar{\nabla}_{\lambda}\bar{\nabla}_{\nu}h_{\mu}^{\penalty 10000\ \lambda}-\bar{\Box}h_{\mu\nu}\right]$
$R_{\mu\nu}^{\text{\tiny{(2)}}}=\frac{\kappa^{2}}{2}\Big[\bar{\nabla}_{\sigma}\left[h^{\sigma\rho}\bar{\nabla}_{\rho}h_{\mu\nu}\right]-\bar{\nabla}_{\sigma}h_{\mu}^{\penalty 10000\ \rho}\bar{\nabla}_{\rho}h_{\nu}^{\penalty 10000\ \sigma}+\bar{\nabla}^{\rho}h_{\mu}^{\penalty 10000\ \sigma}\bar{\nabla}_{\rho}h_{\nu\sigma}-\frac{1}{2}\bar{\nabla}_{\mu}h_{\rho\sigma}\bar{\nabla}_{\nu}h^{\rho\sigma}\Big]$
$R_{\mu\nu}^{\text{\tiny{(3)}}}=-\frac{\kappa^{3}}{2}h^{\rho\sigma}\bar{\nabla}_{\rho}h_{\mu\lambda}\bar{\nabla}_{\sigma}h_{\nu}^{\penalty 10000\ \lambda}$ (11)
The first term is essentially the Fierz-Pauli [5] equation for traceless spin 2 particles in flat space. Hereinafter we will use
$FP_{\mu\nu}\equiv\frac{1}{2}\left[\bar{\nabla}_{\lambda}\bar{\nabla}_{\mu}h_{\nu}^{\penalty 10000\ \lambda}+\bar{\nabla}_{\lambda}\bar{\nabla}_{\nu}h_{\mu}^{\penalty 10000\ \lambda}-\bar{\Box}h_{\mu\nu}\right]$ (12)
in the references it was proved that
$\{R_{\mu\nu}^{\text{\tiny{(1)}}}=0\}\Longrightarrow\{R_{\mu\nu}^{\text{\tiny{(2)}}}+R_{\mu\nu}^{\text{\tiny{(3)}}}=0\}$ (13)
This means that when the deformation $h_{\mu\nu}$ obeys the Fierz-Pauli equation [5] (please remember that $h^{\lambda}_{\lambda}=0$) the deformed metric is Ricci flat.
$R_{\mu\nu}^{\text{\tiny{(1)}}}=\kappa FP_{\mu\nu}=0$ (14)
implies
$R_{\mu\nu}=\bar{R}_{\mu\nu}$ (15)
In the next section, we will present a completely different alternative.
2 The non null Kerr-Schild gauge
In the present work we would like to generalize a little bit this setup by assuming a finite expansion for the inverse metric in terms of a parameter $\xi$
$g^{\mu\nu}=\bar{g}^{\mu\nu}+\xi\kappa h^{\mu\nu}$ (16)
where $\xi$ is an arbitrary constant. In fact we will be able to get $g_{\mu\lambda}g^{\nu\lambda}=\delta_{\mu}^{\nu}$ provided
$\kappa h_{\mu\alpha}\bar{g}^{\alpha\beta}h_{\beta\nu}=-\left(1+\frac{1}{\xi}\right)h_{\mu\nu}$
$\kappa\bar{g}^{\alpha\beta}h_{\alpha\beta}=-\left(1+\frac{1}{\xi}\right)$ (17)
On the other hand, the determinant $g$, using the matrix determinant lemma [7, 6], is related to the background determinant $\bar{g}$
$g=\bar{g}(1+\kappa\bar{g}^{\mu\nu}h_{\mu\nu})=-\frac{1}{\xi}\bar{g}$ (18)
note, if $\xi>0$, we can change the signature of the metric.
It is easy to check that there is also a relationship between the determinant of the deformation $h_{\mu\nu}$ and the determinant of the background metric $\bar{g}_{\mu\nu}$,
$\kappa^{4}h^{2}=\left(1+\frac{1}{\xi}\right)^{4}\bar{g}h$ (19)
There are two solutions. Either $h=0$, or else
$h=\frac{1}{\kappa^{4}}\left(1+\frac{1}{\xi}\right)^{4}\,\bar{g}$ (20)
but this second possibility corresponds to a rescaling of the metric
$h_{\mu\nu}=-\frac{1}{\kappa}\left(1+\frac{1}{\xi}\right)\bar{g}_{\mu\nu}$ (21)
and the total metric is
$g_{\mu\nu}=-\frac{1}{\xi}\bar{g}_{\mu\nu}$ (22)
Forgetting about this trivial rescaling, we will concentrate on the case $h=0$. Ttis means that the matrix $h$ has got rank $r\leq 3$. Therefore, the condition (2) has three different solutions
  • Rank one
    $h_{\mu\nu}=\kappa A_{\mu}A_{\nu}$ (23)
    there is a three-dimensional subspace of zero modes, defined by the condition
    $A_{\mu}n_{i}^{\mu}=0\quad i=1,2,3$ (24)
    the condition (2) is fulfilled as long as
    $\bar{g}^{\alpha\beta}A_{\alpha}A_{\beta}=A^{\lambda}A_{\lambda}=-\frac{1}{\kappa^{2}}\left(1+\frac{1}{\xi}\right)$ (25)
    in conclusion the metric reads
    $g_{\mu\nu}=\bar{g}_{\mu\nu}+\kappa^{2}A_{\mu}A_{\nu}$ (26)
    and the inverse of the metric
    $g^{\mu\nu}=\bar{g}^{\mu\nu}+\xi\kappa^{2}A^{\mu}A^{\nu}$ (27)
    in particular the determinant
    $g=\left(1+\kappa^{2}A^{\lambda}A_{\lambda}\right)\bar{g}$ (28)
    This shows that for $\kappa^{2}A^{2}+1>0$ (that is, for timelike or null vector fields, and even for spacelike ones with $\kappa^{2}A^{2}>-1$) both $g$ and $\bar{g}$ have got the same sign; whereas for $\kappa^{2}A^{2}-1<0$ they have opposite sign: there is a change from a Euclidean (Riemannian) metric to a Lorenztian (Pseudo-Riemannian) one.
  • Rank two. Higher rank is more restrictive.
    $h_{\mu\nu}=\kappa A_{\mu}A_{\nu}+\kappa B_{\mu}B_{\nu}$ (29)
    with
    $A^{\lambda}A_{\lambda}=B^{\lambda}B_{\lambda}=-\frac{1}{2\kappa^{2}}\left(1+\frac{1}{\xi}\right)$
    $A^{\lambda}B_{\lambda}=0$ (30)
    this means that the vectors AA and $B$ are orthogonal and equally normalized. They cannot be both null. This condition could be softened by allowing two different $\xi_{A}\neq\xi_{B}$. There is a two-dimensional subspace of zero modes
    $A_{\mu}n_{i}^{\mu}=B_{\mu}n_{i}^{\mu}=0\quad i=1,2$ (31)
  • Rank three
    $h_{\mu\nu}=\kappa A_{\mu}A_{\nu}+\kappa B_{\mu}B_{\nu}+\kappa C_{\mu}C_{\nu}$ (32)
    where the vectors A, B and C are independent and non-null, in fact
    $A^{\lambda}A_{\lambda}=B^{\lambda}B_{\lambda}=C^{\lambda}C_{\lambda}=-\frac{1}{3\kappa^{2}}\left(1+\frac{1}{\xi}\right)$
    $A^{\lambda}B_{\lambda}=A^{\lambda}C_{\lambda}=B^{\lambda}C_{\lambda}=0$ (33)
    and there is now only a one-dimensional subspace of zero modes
    $A_{\mu}n^{\mu}=B_{\mu}n^{\mu}=C_{\mu}n^{\mu}=0$ (34)
From now on, we shall concentrate for simplicity in the rank one case. We will study the expression of the Ricci tensor, and its properties when applying the condition (25), concluding with a illustrative example. To begin with, the connection of the metric (16) is
$\Gamma^{\lambda}_{\mu\nu}=\bar{\Gamma}^{\lambda}_{\mu\nu}+C^{\lambda}_{\mu\nu}$ (35)
where
$C^{\lambda}_{\mu\nu}=\frac{\kappa^{2}}{2}\left[A_{\mu}\bar{\nabla}_{\nu}A^{\lambda}-\xi A^{\lambda}\bar{\nabla}_{\nu}A_{\mu}+A_{\nu}\bar{\nabla}_{\mu}A^{\lambda}-\xi A^{\lambda}\bar{\nabla}_{\mu}A_{\nu}-\bar{\nabla}^{\lambda}(A_{\mu}A_{\nu})-\xi\kappa^{2}A^{\lambda}A^{\tau}\bar{\nabla}_{\tau}(A_{\mu}A_{\nu})\right]$
with $C^{\lambda}_{\lambda\mu}=0$ The Ricci tensor of the deformed metric is
$R_{\mu\nu}=\bar{R}_{\mu\nu}+R_{\mu\nu}^{\text{\tiny{(1)}}}+R_{\mu\nu}^{\text{\tiny{(2)}}}+R_{\mu\nu}^{\text{\tiny{(3)}}}$ (37)
where to be specific,
$R_{\mu\nu}^{\text{\tiny{(1)}}}$ $=\frac{\kappa^{2}}{2}\Big[FP_{\mu\nu}-(1+\xi)\left[\bar{\nabla}_{\lambda}[A^{\lambda}\bar{\nabla}_{\mu}A_{\nu}]+\bar{\nabla}_{\lambda}[A^{\lambda}\bar{\nabla}_{\nu}A_{\mu}]-\bar{\nabla}_{\lambda}A_{\mu}\bar{\nabla}^{\lambda}A_{\nu}+\bar{\nabla}_{\mu}A^{\lambda}\bar{\nabla}_{\nu}A_{\lambda}\right]\Big]$
$R_{\mu\nu}^{\text{\tiny{(2)}}}$ $=\frac{\kappa^{4}}{2}\Big[A_{\mu}A_{\nu}\left(\bar{\nabla}_{\alpha}A_{\beta}\bar{\nabla}^{\alpha}A^{\beta}-\bar{\nabla}_{\alpha}A_{\beta}\bar{\nabla}^{\beta}A^{\alpha}\right)-\xi A^{\alpha}\bar{\nabla}_{\alpha}(A_{\mu}A_{\nu})\bar{\nabla}_{\beta}A^{\beta}$-
$-\xi A^{\alpha}A^{\beta}\left(A_{\mu}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}A_{\nu}+A_{\nu}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}A_{\mu}+\bar{\nabla}_{\alpha}A_{\mu}\bar{\nabla}_{\beta}A_{\nu}\right)\Big]$
$R_{\mu\nu}^{\text{\tiny{(3)}}}$ $=\xi\frac{\kappa^{6}}{2}A_{\mu}A_{\nu}A^{\alpha}A^{\beta}\bar{\nabla}_{\alpha}A^{\lambda}\bar{\nabla}_{\beta}A_{\lambda}$ (38)
It is easy to see that now, as opposed to what happens in the null case, FP is not enough to map Ricci flat spaces into Ricci flat spaces.
$\{R_{\mu\nu}^{\text{\tiny{(1)}}}=0\}\nRightarrow\{R_{\mu\nu}^{\text{\tiny{(2)}}}+R_{\mu\nu}^{\text{\tiny{(3)}}}=0\}$ (39)
then in general
$\{R_{\mu\nu}=0\}\neq\{\bar{R}_{\mu\nu}=0\}$ (40)
We shall present now an explicit example of this fact
Start with a metric of Euclidean signature
$\text{d}s^{2}=-2\text{d}t\text{d}z-2\text{d}x\text{d}y$ (41)
with $\bar{R}_{\mu\nu}=0$. Add the deformation
$h_{\mu\nu}=\begin{pmatrix}\frac{(\chi-2C_{1}h_{2}[t])^{2}}{4C^{2}_{3}}&\frac{C_{1}(\chi-2C_{1}h_{2}[t])}{2C_{3}}&\frac{h_{2}[t](\chi-2C_{1}h_{2}[t])}{2C_{3}}&\frac{1}{2}\left(\chi-2C_{1}h_{2}[t]\right)\\ \frac{C_{1}(\chi-2C_{1}h_{2}[t])}{2C_{3}}&C_{1}^{2}&C_{1}h_{2}[t]&C_{1}C_{3}\\ \frac{h_{2}[t](\chi-2C_{1}h_{2}[t])}{2C_{3}}&C_{1}h_{2}[t]&h_{2}^{2}[t]&h_{2}[t]C_{3}\\ \frac{1}{2}\left(\chi-C_{1}h_{2}[t]\right)&C_{1}C_{3}&h_{2}[t]C_{3}&C_{3}^{2}\end{pmatrix}$ (42)
where the function $h_{2}[t]$ has the property $h_{2}^{\prime\prime}[t]=0$ and $\chi\equiv 1+\frac{1}{\xi}$. Then this deformation $h_{\mu\nu}$ obeys the Fierz-Pauli equation
$FP_{\mu\nu}=0$ (43)
The total metric, with signature $g=1-\chi$
$\text{d}s^{2}=-2\text{d}t\text{d}z-2\text{d}x\text{d}y+\left[\frac{\chi-2C_{1}h_{2}[t]}{2C_{3}}\text{d}t+C_{1}\text{d}x+h_{2}[t]\text{d}y+C_{3}\text{d}z\right]^{2}$ (44)
is not Ricci flat , however
$R_{\mu\nu}=\frac{(h_{2}^{\prime}[t])^{2}}{2-2\chi}\begin{pmatrix}C_{1}^{2}&0&-C_{1}C_{3}&0\\ 0&0&0&0\\ -C_{1}C_{3}&0&C_{3}^{2}&0\\ 0&0&0&0\end{pmatrix}$ (45)
A natural question to ask at this moment is the following. What are the conditions that the deformation must obey (we already know that it is not true in general) in order to keep Ricci flatness? In particular this would hold true whenever
$R_{\mu\nu}=\bar{R}_{\mu\nu}$ (46)
We will answer this question in the next section.
3 The general deformation that obeys $R_{\mu\nu}=\bar{R}_{\mu\nu}$
In this section we want to obtain the necessary conditions that a deformation must fulfill in order for it to map vacuum solutions into vacuum solutions. Let us start with the necessary but not sufficient condition that the scalar curvature be invariant $R=\bar{R}$. In fact,
$R=\bar{R}-\kappa^{2}\xi\left[\bar{\nabla}_{\alpha}A^{\alpha}\bar{\nabla}_{\beta}A^{\beta}+A^{\alpha}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}A^{\beta}+A^{\alpha}\bar{\nabla}_{\beta}\bar{\nabla}_{\alpha}A^{\beta}-A^{\alpha}\bar{\nabla}_{\beta}\bar{\nabla}^{\beta}A^{\alpha}\right]+$
$+\frac{\kappa^{2}}{2}(-1+\xi)\left[\bar{\nabla}_{\alpha}A_{\beta}\bar{\nabla}^{\alpha}A^{\beta}-\bar{\nabla}_{\alpha}A_{\beta}\bar{\nabla}^{\beta}A^{\alpha}\right]-\xi\frac{\kappa^{4}}{2}A^{\mu}A^{\nu}\bar{\nabla}_{\mu}A_{\lambda}\bar{\nabla}_{\nu}A^{\lambda}$ (47)
Assume that the perturbation satisfies the Fierz-Pauli equation
$FP_{\mu\nu}\equiv\frac{1}{2}\left[\bar{\nabla}_{\lambda}\bar{\nabla}_{\mu}(A_{\nu}A^{\lambda})+\bar{\nabla}_{\lambda}\bar{\nabla}_{\nu}(A_{\mu}A^{\lambda})-\bar{\Box}(A_{\mu}A_{\nu})\right]=0$ (48)
Multiply by $A_{\mu}A_{\nu}$
$A_{\alpha}\bar{\nabla}_{\beta}\bar{\nabla}^{\beta}A^{\alpha}-A_{\alpha}\bar{\nabla}_{\beta}\bar{\nabla}^{\alpha}A^{\beta}=\frac{\xi}{1+\xi}\kappa^{2}A^{\mu}A^{\nu}\bar{\nabla}_{\mu}A_{\lambda}\bar{\nabla}_{\nu}A^{\lambda}$ (49)
Taking the trace, that is, multiplying by $\bar{g}^{\mu\nu}$
$\bar{\nabla}_{\alpha}A^{\alpha}\bar{\nabla}_{\beta}A^{\beta}+A^{\alpha}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}A^{\beta}+A^{\alpha}\bar{\nabla}_{\beta}\bar{\nabla}_{\alpha}A^{\beta}-A_{\alpha}\bar{\nabla}_{\beta}\bar{\nabla}^{\beta}A^{\alpha}+\bar{\nabla}_{\alpha}A_{\beta}\bar{\nabla}^{\beta}A^{\alpha}-\bar{\nabla}_{\alpha}A_{\beta}\bar{\nabla}^{\alpha}A^{\beta}=0$ (50)
Plugging inside the scalar of curvature both results we learn that
$R=\bar{R}+\frac{\kappa^{2}}{2}(1+\xi)\bar{\nabla}_{\alpha}\left[A^{\beta}\bar{\nabla}_{\beta}A^{\alpha}\right]$ (51)
This means that to have $R=\bar{R}$, we have three options, either $\xi=-1$, which corresponds with the null Kerr-Schild or else
$A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=0$ (52)
or the third possibility,
$A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=\phi A_{\beta}\hskip 28.45274pt\text{with}\hskip 28.45274pt\bar{\nabla}_{\alpha}A^{\alpha}=0$ (53)
Let us study the new conditions in turn.
3.1 The first condition $A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=0$
To begin with, automatically $R_{\mu\nu}^{\text{\tiny{(3)}}}=0$, and our previous analysis of the FP equation imply that
$\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\beta}A^{\alpha}-\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\alpha}A^{\beta}=0\Longrightarrow\bar{\nabla}_{\mu}A_{\nu}=\bar{\nabla}_{\nu}A_{\mu}$ (54)
This leads to
$R_{\mu\nu}^{\text{\tiny{(1)}}}=\frac{\kappa^{2}}{2}\Big[FP_{\mu\nu}-(1+\xi)\left[\bar{\nabla}_{\lambda}[A^{\lambda}\bar{\nabla}_{\mu}A_{\nu}]+\bar{\nabla}_{\lambda}[A^{\lambda}\bar{\nabla}_{\nu}A_{\mu}]\right]\Big]=-\xi\frac{\kappa^{2}}{2}FP_{\mu\nu}=0\\$
as well as
$R_{\mu\nu}^{\text{\tiny{(2)}}}=-\xi\frac{\kappa^{4}}{2}A^{\alpha}A^{\beta}\left(A_{\mu}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}A_{\nu}+A_{\nu}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}A_{\mu}\right)=-\xi\frac{\kappa^{4}}{2}A^{\alpha}A^{\beta}\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}(A_{\mu}A_{\nu})=0\\$
to conclude
$FP_{\mu\nu}=0\hskip 5.69046pt\text{with}\hskip 5.69046ptA^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=0\Rightarrow R_{\mu\nu}=\bar{R}_{\mu\nu}$ (55)
This means that whenever the deformation is geodesic, it maintains Ricci flatness.
3.2 The second alternative $A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=\phi A_{\beta}$ and $\bar{\nabla}_{\mu}A^{\mu}=0$
In this case,(49), implies
$\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\beta}A^{\alpha}-\bar{\nabla}_{\beta}A^{\alpha}\bar{\nabla}_{\alpha}A^{\beta}=\phi^{2}$ (56)
and a bit of algebra leads to
$R_{\mu\nu}=\bar{R}_{\mu\nu}-(1+\xi)\frac{\kappa^{2}}{2}\left[A^{\lambda}\bar{\nabla}_{\lambda}\left(\bar{\nabla}_{\mu}A_{\nu}+\bar{\nabla}_{\nu}A_{\mu}\right)]-\bar{\nabla}_{\lambda}A_{\mu}\bar{\nabla}^{\lambda}A_{\nu}+\bar{\nabla}_{\mu}A^{\lambda}\bar{\nabla}_{\nu}A_{\lambda}\right]-\xi\kappa^{4}\phi^{2}A_{\mu}A_{\nu}$
in conclusion, we get the desired result $R_{\mu\nu}=\bar{R}_{\mu\nu}$ only when $\phi=0$, in which case this case the second condition reduces to the previous one.
4 Conclusions
We have generalized the usual null Kerr-Schild [3] gauge
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\kappa^{2}A_{\mu}A_{\nu}$ (58)
with $A^{2}=0$, to the non null case, allowing any value for $A^{2}$. This needs to assume that the inverse metric is given by
$g^{\mu\nu}=\bar{g}^{\mu\nu}+\xi\kappa^{2}A^{\mu}A^{\nu}$ (59)
where
$A^{\lambda}A_{\lambda}=-\frac{1}{\kappa^{2}}\left(1+\frac{1}{\xi}\right)$ (60)
The curvature tensor is now given by a finite sum of terms, but in contraposition to the null case, however, the said Fierz-Pauli equation is not enough to map Ricci flat spaces amongst themselves.
We have been able to prove a theorem however, that states that when the deformation is geodesic of the background metric, that is when
$A^{\lambda}\bar{\nabla}_{\lambda}A_{\sigma}=0$ (61)
then the Ricci tensor is fully determined by its linear piece, so that if the deformation obeys the the Fierz-Pauli equation then the Ricci tensor is invariant. This obviously means that whenever the background spacetime is Ricci flat, so is the deformed one.
$\bar{R}_{\mu\nu}=R_{\mu\nu}$ (62)
5 Acknowledgements
We acknowledge partial financial support by the Spanish MINECO through the Centro de excelencia Severo Ochoa Program under Grant CEX2020-001007-S funded by MCIN/AEI/10.13039/501100011033. We also acknowledge partial financial support by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant PID2022-137127NB-I00 funded by MCIN/AEI/10.13039/501100011033/ FEDER, UE. All authors acknowledge the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 860881-HIDDeN and also byGrant PID2019-108892RB-I00 funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF A way of making Europe”.